# GraphThinker: Reinforcing Video Reasoning with Event Graph Thinking

### v1

@@ 1 Introduction
Figure 1: Current MLLMs [2] implicitly model event relations in video by token correlations. This often leads to hallucinations in video reasoning. For instance, when determining the temporal order of events, QwenVL-2.5 tends to yield temporal hallucinations due to a lack of explicit event-level evidence validation and reasoning. In contrast, our GraphThinker explicitly models both intra- and inter-event relations in accordance with selective visual attentions in a video to constrain reasoning under these structured event relations anchored by fine-grained visual evidence, reducing hallucinations and improving temporal consistency in reasoning.
@@ 4.1 Experimental Setup
QwenVL-2.5 [2] 7B 50.46 81.99 67.50 46.77
QwenVL-2.5+EVSG 7B 50.92 83.23 75.33 54.58
@@
Implementation Details. We used QwenVL-2.5-7B [2] as our baseline due to its strong video understanding ability and open-source availability. We also use this baseline for EVSG generation, including generating the multi-grained dense captions and building the EVSG process. The event number limit for dense caption is set to (5, 10, 15). We set $\lambda_{1}{=}0.7$, $\lambda_{2}{=}0.3$, $\lambda_{3}{=}0.6$ and $\alpha{=}0.3$ for the reward function of GRPO post-training. We also utilise the same baseline for event graph generation and input the generated graph during inference. All experiments are conducted using PyTorch on 8 NVIDIA A100 GPUs, with a maximum prompt length of 4096, response length of 2048, 8 rollouts, a batch size of 16, and 1 training epoch.
@@ 4.2 Comparisons of the state-of-the-arts
VideoHalluc. Tab. 2 shows that our method effectively mitigates the video hallucination problems for the baseline model. When integrating the EVSG to QwenVL-2.5 without any post-training, GraphThinker’s performance across all tasks in the VidHalluc benchmark is notably improved over baseline, especially in the Temporal Sequence Hallucination (TSH) and Scene Transition Hallucination (STH) tasks, improving 7.83% and 7.81%, respectively. These results demonstrate that EVSG’s explicit modeling of event relations in video helps better model understanding in reasoning compared to the implicit event correlation in the baseline. Moreover, GraphThinker also has slight improvement in better action recognition. After conducting GRPO finetuning on Rextime for leveraging EVSG better, GraphThinker achieves the state-of-the-art across the open-source MLLMs. Closed-source models GPT-4o and Gemini-1.5-Pro outperform all open-source models in Tab. 2, partly due to their much larger sizes (GPT-4o: 200B; Gemini-1.5-Pro: 200–600B parameters depending on MoE configuration).
Figure 4: A visual example showing that our method reduces hallucination during reasoning. QwenVL-2.5 still yields hallucination in reasoning during the inference stage. Our method explicitly models event relations to guide the reasoning process for MLLM, producing visually grounded and temporally consistent reasoning.
@@
We conducted ablation studies with the QwenVL-2.5-3B model on the validation set of the Rextime benchmark to systematically evaluate the contribution of each component in our approach. In Tab. 4.3, we first conduct training-free testing in row 1-3. The results show that our EVSG can benefit the model to understand the event causal relation and achieve more consistent semantic understanding and temporal localization, improving 7.93% on Acc@0.5. Besides, our EVSG demonstrates a better event relation modeling than dense captioning (DC), higher than 3.48% on Acc@0.5. Then, we evaluate post-training testing. Applying GRPO alone yields moderate improvements, but combining it with our EVSG leads to significant gains across all metrics. Notably, the integration of the $r_{attn}$ in GRPO further achieves better performance, showing that our proposed $r_{attn}$ better leverages the EVSG and high-value visual cues in videos for reasoning.
@@
QwenVL-2.5-7B 13.77 55.65 6.20
QwenVL-2.5-7B + DC 24.25 62.87 15.42
QwenVL-2.5-7B + EVSG 25.36 63.30 16.07
@@
QwenVL-2.5-7B 41.04 57.55 27.03
QwenVL-2.5-7B + DC 41.20 69.60 29.10
QwenVL-2.5-7B + DC + $r_{attn}$ 41.42 71.01 32.72
QwenVL-2.5-7B + EVSG 42.57 72.20 32.03
QwenVL-2.5-7B + EVSG + $r_{attn}$ 43.56 73.72 33.49
Table 3: Ablation study of QwenVL-2.5-7B with various training strategies on the validation set of the Rextime benchmark. “DC” denotes dense captioning.
@@
Table 4: Parameter analysis of QwenVL-2.5-7B+EVSG under the training-free setting on the Rextime validation set. The event number limit controls the maximum number of events generated in dense captioning for event graph generation.

### v2

@@ 1 Introduction
Figure 1: Current MLLMs [2] implicitly model event relations in video by token correlations. This often leads to hallucinations in video reasoning. For instance, when determining the temporal order of events, Qwen2.5-VL tends to yield temporal hallucinations due to a lack of explicit event-level evidence validation and reasoning. In contrast, our GraphThinker explicitly models both intra- and inter-event relations in accordance with selective visual attentions in a video to constrain reasoning under these structured event relations anchored by fine-grained visual evidence, reducing hallucinations and improving temporal consistency in reasoning.
@@ 4.1 Experimental Setup
Qwen2.5-VL [2] 7B 50.46 81.99 67.50 46.77
Qwen2.5-VL+EVSG 7B 50.92 83.23 75.33 54.58
@@
Implementation Details. We used Qwen2.5-VL-7B [2] as our baseline due to its strong video understanding ability and open-source availability. We also use this baseline for EVSG generation, including generating the multi-grained dense captions and building the EVSG process. The event number limit for dense caption is set to (5, 10, 15). We set $\lambda_{1}{=}0.7$, $\lambda_{2}{=}0.3$, $\lambda_{3}{=}0.6$ and $\alpha{=}0.3$ for the reward function of GRPO post-training. We also utilise the same baseline for event graph generation and input the generated graph during inference. All experiments are conducted using PyTorch on 8 NVIDIA A100 GPUs, with a maximum prompt length of 4096, response length of 2048, 8 rollouts, a batch size of 16, and 1 training epoch.
@@ 4.2 Comparisons of the state-of-the-arts
VideoHalluc. Tab. 2 shows that our method effectively mitigates the video hallucination problems for the baseline model. When integrating the EVSG to Qwen2.5-VL without any post-training, GraphThinker’s performance across all tasks in the VidHalluc benchmark is notably improved over baseline, especially in the Temporal Sequence Hallucination (TSH) and Scene Transition Hallucination (STH) tasks, improving 7.83% and 7.81%, respectively. These results demonstrate that EVSG’s explicit modeling of event relations in video helps better model understanding in reasoning compared to the implicit event correlation in the baseline. Moreover, GraphThinker also has slight improvement in better action recognition. After conducting GRPO finetuning on Rextime for leveraging EVSG better, GraphThinker achieves the state-of-the-art across the open-source MLLMs. Closed-source models GPT-4o and Gemini-1.5-Pro outperform all open-source models in Tab. 2, partly due to their much larger sizes (GPT-4o: 200B; Gemini-1.5-Pro: 200–600B parameters depending on MoE configuration).
Figure 4: A visual example showing that our method reduces hallucination during reasoning. Qwen2.5-VL still yields hallucination in reasoning during the inference stage. Our method explicitly models event relations to guide the reasoning process for MLLM, producing visually grounded and temporally consistent reasoning.
@@
We conducted ablation studies with the Qwen2.5-VL-3B model on the validation set of the Rextime benchmark to systematically evaluate the contribution of each component in our approach. In Tab. 4.3, we first conduct training-free testing in row 1-3. The results show that our EVSG can benefit the model to understand the event causal relation and achieve more consistent semantic understanding and temporal localization, improving 7.93% on Acc@0.5. Besides, our EVSG demonstrates a better event relation modeling than dense captioning (DC), higher than 3.48% on Acc@0.5. Then, we evaluate post-training testing. Applying GRPO alone yields moderate improvements, but combining it with our EVSG leads to significant gains across all metrics. Notably, the integration of the $r_{attn}$ in GRPO further achieves better performance, showing that our proposed $r_{attn}$ better leverages the EVSG and high-value visual cues in videos for reasoning.
@@
Qwen2.5-VL-7B 13.77 55.65 6.20
Qwen2.5-VL-7B + DC 24.25 62.87 15.42
Qwen2.5-VL-7B + EVSG 25.36 63.30 16.07
@@
Qwen2.5-VL-7B 41.04 57.55 27.03
Qwen2.5-VL-7B + DC 41.20 69.60 29.10
Qwen2.5-VL-7B + DC + $r_{attn}$ 41.42 71.01 32.72
Qwen2.5-VL-7B + EVSG 42.57 72.20 32.03
Qwen2.5-VL-7B + EVSG + $r_{attn}$ 43.56 73.72 33.49
Table 3: Ablation study of Qwen2.5-VL-7B with various training strategies on the validation set of the Rextime benchmark. “DC” denotes dense captioning.
@@
Table 4: Parameter analysis of Qwen2.5-VL-7B+EVSG under the training-free setting on the Rextime validation set. The event number limit controls the maximum number of events generated in dense captioning for event graph generation.